# 3D Software Synthesis Guided by Constraint-Expressive Intermediate Representation

### v1

@@ 1. Introduction
Software Root (Scenes)Asset (Object) AAsset (Object) BMeshRendererTransformMaterialMeshRendererTransformScriptVerticesFacesUV MapsPosition$(x,y,z)$Rotation$(r_{x},r_{y},r_{z})$Scale$(s_{x},s_{y},s_{z})$Typical 3D Software Hierarchy
@@ 2. Preliminaries
Spatial Constraints. Professional 3D software must satisfy three constraint categories: Geometric constraints specify spatial relationships (e.g., ”A is 2m from B”: $||pos_{A}-pos_{B}||_{2}=2.0$). Physical constraints ensure plausibility through collision avoidance and gravity support. Semantic constraints encode domain rules (e.g., ”emergency exits must be accessible”).
@@ 3.1. Overview and Design Principles
[…]
@@
[…]
@@ 3.2.1. Natural Language Analysis and Contextualization
Scenethesis then performs controlled prompt expansion based on LLMs to enrich the description with contextual details, since user requirements usually contain hidden constraints. For example, the user requirement “a modern conference room” contains hidden requirements for the furniture arrangements, lighting conditions, and accessibility. The expansion is strictly constrained to preserve all explicit user requirements while adding plausible hidden constraints inferred from the user requirements. Formally, let $Q^{{}^{\prime}}$ denote the expanded prompt where $Q^{{}^{\prime}}=Q\cup\{c_{1},c_{2},\ldots,c_{k}\}$ such that each $c_{i}$ represents an inferred contextual constraint from $Q$. Next, sub-prompt $Q^{{}^{\prime}}_{i}$ ($i\in\left\{1,\dots,r\right\}$ where $r$ is the number of regions) for region $i$ is generated via another LLM call by extracting sentences in $Q^{{}^{\prime}}$ that are relevant to region $i$. In the following stages, scene-wise processing and generation use $Q^{{}^{\prime}}$ while region-wise processing and generation use $Q^{{}^{\prime}}_{i}$.
@@ 3.2.2. DSL Specification Generation
[…]
@@
3.3. Stage II: Asset Synthesis
@@ 3.4.1. Iterative Constraint Resolution Algorithm
3:$L_{0}\leftarrow\textsc{InitialPlacement}(\mathcal{O})$
4:$L_{0}\leftarrow\textsc{PhysicsRelaxation}(L_{0})$
@@
The core solving loop (lines 3$\sim$10) iteratively addresses unsatisfied constraints in batches. For each batch $\mathcal{B}$ of size $k$, we invoke LLMSolve, which leverages the spatial reasoning capabilities of LLMs to suggest object transformations. The LLM receives a structured description of the current layout, the violated constraints, and the complete constraint set, then proposes specific adjustments (translations, rotations) to resolve the violations.
@@ 5.4.1. RQ1 and RQ2
For generating our dataset, we use deepseek-v3-0324 (Liu et al., 2024). For running Scenethesis and baselines, we use gpt-4o-2024-11-20 (Hurst et al., 2024), gemini-2.5-pro-preview-06-05 (Comanici et al., 2025), and deepseek-r1-250528 (Guo et al., 2025). For object canonical orientation detection and Visual Question Answering (VQA, one of our visual metrics), we use claude-3-7-sonnet-20250219 (Anthropic, 2025). The non-repetitive use of LLMs ensures that experimental results are less likely to be biased towards a particular LLM backbone. Though, it should be noted that Scenethesis supports any LLM with sufficient reasoning capabilities. Also, the temperature parameter for all LLM calls is set to $0.7$ (except for the use of 00 in scene type classification at Stage I, object canonical orientation detection at Stage II, and VQA).
@@ 5.5.1. Stage I: Constraint Resemblance
In ScenethesisLang, constraints are divided into two types: object constraints and layout constraints. To evaluate whether Scenethesis can generate object constraints that match those in the ground truth (i.e., our dataset), we first use Phrase-BERT (Wang et al., 2021a) and Sentence-BERT (Reimers and Gurevych, 2019) to compute the high-dimensional embeddings for object names and descriptions respectively. Then, we compute the dot-product (scaled to $\left[0,1\right]$) between every pair of object names as well as every pair of object descriptions. The confidence that a generated object matches a ground truth object is the harmonic mean between the corresponding “name” and “description” scaled dot-products. Next, we use the Hungarian algorithm to create a one-to-one mapping from the generated objects to the ground truth objects. Entries in this matrix that are smaller than threshold $\tau_{o}$ are zeroed out. Finally, to compute the F1 score ($\frac{2\times\text{precision}\times\text{recall}}{\text{precision}+\text{recall}}$, where $\text{precision}=\frac{TP}{TP+FP}$ and $\text{recall}=\frac{TP}{TP+FN}$), we define $TP$ as the number generated objects that are mapped to one (and only one) ground truth object, $FP$ as the number of generated objects that are not mapped to any ground truth object, and $FN$ as the number of ground truth objects that are not mapped to any generated object.
As for layout constraints, we first translate each generated constraint into some NL (i.e., an intuitive and human-understandable sentence). Then, we use Sentence-BERT to compute the embeddings of the translated generated constraints and ground truth NL-driven constraints. We then compute the scaled dot-product between every pair of generated and ground truth constraints, creating a confidence matrix (in some sense, a many-to-many mapping). For each entry in this matrix, if either (1) no object name from the corresponding ground truth constraint exists in the corresponding generated constraint or (2) it is smaller than a threshold $\tau_{l}$, the entry is zeroed out. Finally, to compute the F1 score, we define $TP$ as the number of ground truth constraints that are mapped to at least one generated constraint, $FP$ as the number of unmapped generated constraints, and $FN$ as the number of unmapped ground truth constraints.
@@ 5.5.2. Stage II: Object-Query Coherence
For each acquired object, we (1) compute its smallest bounding sphere (with radius $r$ units), (2) place a camera $r\sin{\frac{FOV}{2}}$ units away from the object (where $FOV$ (in radian) is the field of view of the camera) pointing towards the front-facing side of the object, and (3) render the camera view using Blender on a white background. We then use BLIP-2 (Li et al., 2023a) (with its ITM head) and CLIP (Radford et al., 2021; Hessel et al., 2021) to measure the coherence between the formulated object query generated by Scenethesis and the rendered image. To reduce bias, in addition to the object query itself, we further pass a combination of “a 3D model of ” followed by the object query to the evaluation tools, and the final tool-wise score is the maximum between the two trials. The overall coherence is the arithmetic mean between the final BLIP and CLIP scores.

### v2

@@ 1. Introduction
[…]
@@ 2. Preliminaries
Spatial Constraints. Professional 3D software must satisfy three constraint categories: Geometric constraints specify spatial relationships (e.g., ”A is 2m from B”: $pos_{A}-pos_{B}||_{2}=2.0$). Physical constraints ensure plausibility through collision avoidance and gravity support. Semantic constraints encode domain rules (e.g., ”emergency exits must be accessible”).
@@ 3.1. Overview and Design Principles
(1) $program\in\mathit{Programs}::=$ $\ stmt;\ |\ stmt;\ program$
(2) $stmt\in\mathit{Statements}::=$ $\ \mathit{decl}\ |\ \mathit{const}\ |\ assign$
(3) $decl\in\mathit{Declarations}::=$ $\ \textbf{object}\;id\ |\ \textbf{region}\;id\ |\ \tau\ id$
(4) $const\in\mathit{Constraints}::=$ $\ \textbf{assert}\;\phi\ |\ \textbf{allowCollide}(id,id)\ |\ \textbf{allowOutside}(id)$
(5) $assign\in\mathit{Assignments}::=$ $\ id.\alpha\leftarrow e\ |\ id.\beta\leftarrow e\ |\ id\leftarrow e$
(6) $\alpha\in\mathit{Object\ Properties}::=$ $\ \textbf{color}\mid\textbf{material}\mid\textbf{features}$
(7) $\beta\in\mathit{Region\ Properties}::=$ $\ \textbf{pos}\mid\textbf{rot}\mid\textbf{scale}\mid$
(8) $\phi\in\mathit{Assertions}::=$ $\ e\,\bowtie\,e\mid\textbf{inside}(id,id)\mid\phi\land\phi\mid\phi\lor\phi\mid\lnot\phi$
(9) $\tau\in\mathit{Types}::=$ $~\textbf{Number}\mid\textbf{Degree}\mid\textbf{Bool}\mid\textbf{Vector3}$
(10) $\mid\textbf{Rotation}\mid\textbf{Color}\mid\textbf{Material}$
(11) $e\in\mathit{Expressions}::=$ $\ n\mid id\mid s\mid e\,\odot\,e$
@@
(14) $\bowtie\in\mathit{Compare\ Operators}::=$ $~=\mid\neq\mid<\mid\leq\mid>\mid\geq$
(15) $\odot\in\mathit{Arithmetic\ Operators}::=$ $~+\mid-\mid*\mid/$
(16) $n\in\mathit{Numbers}::=$ any number
(17) $s\in\mathit{Strings}::=$ any string
(18) $id\in\mathit{Identifiers}::=$ names of the objects, regions and variables
@@ 3.2.1. Natural Language Analysis and Contextualization
Scenethesis then performs controlled prompt expansion based on LLMs to enrich the description with contextual details, since user requirements usually contain hidden constraints. For example, the user requirement “a modern conference room” contains hidden requirements for the furniture arrangements, lighting conditions, and accessibility. The expansion is strictly constrained to preserve all explicit user requirements while adding plausible hidden constraints inferred from the user requirements. Formally, let $Q^{{}^{\prime}}$ denote the expanded prompt where $Q^{{}^{\prime}}=Q\cup\{c_{1},c_{2},\ldots,c_{k}\}$ such that each $c_{i}$ represents an inferred contextual constraint from $Q$. Next, sub-prompt $Q^{{}^{\prime}}_{i}$ ($i\in\left\{1,\dots,r\right\}$ where $r$ is the number of regions) for region $i$ is generated via another LLM call by extracting sentences in $Q^{{}^{\prime}}$ that are relevant to region ii. In the following stages, scene-wise processing and generation use $Q^{{}^{\prime}}$ while region-wise processing and generation use $Q^{{}^{\prime}}_{i}$.
@@ 3.2.2. DSL Specification Generation
$\texttt{assert}~\textit{lamp}\texttt{.pos.y}>\textit{table}\texttt{.pos.y}+\textit{table}\texttt{.scale.y}$
@@
$3.$ Stage II: Asset Synthesis
@@ 3.4.1. Iterative Constraint Resolution Algorithm
[…]
@@
[…]
@@ 5.4.1. RQ1 and RQ2
For generating our dataset, we use deepseek-v3-0324 (Liu et al., 2024). For running Scenethesis and baselines, we use gpt-4o-2024-11-20 (Hurst et al., 2024), gemini-2.5-pro-preview-06-05 (Comanici et al., 2025), and deepseek-r1-250528 (Guo et al., 2025). For object canonical orientation detection and Visual Question Answering (VQA, one of our visual metrics), we use claude-3-7-sonnet-20250219 (Anthropic, 2025). The non-repetitive use of LLMs ensures that experimental results are less likely to be biased towards a particular LLM backbone. Though, it should be noted that Scenethesis supports any LLM with sufficient reasoning capabilities. Also, the temperature parameter for all LLM calls is set to $0.7$ (except for the use of 0 in scene type classification at Stage I, object canonical orientation detection at Stage II, and VQA).
@@ 5.5.1. Stage I: Constraint Resemblance
In ScenethesisLang, constraints are divided into two types: object constraints and layout constraints. To evaluate whether Scenethesis can generate object constraints that match those in the ground truth (i.e., our dataset), we first use Phrase-BERT (Wang et al., 2021a) and Sentence-BERT (Reimers and Gurevych, 2019) to compute the high-dimensional embeddings for object names and descriptions respectively. Then, we compute the dot-product (scaled to $\left[0,1\right]$) between every pair of object names as well as every pair of object descriptions. The confidence that a generated object matches a ground truth object is the harmonic mean between the corresponding “name” and “description” scaled dot-products. Next, we use the Hungarian algorithm to create a one-to-one mapping from the generated objects to the ground truth objects. Entries in this matrix that are smaller than threshold $\tau_{o}$ are zeroed out. Finally, to compute the F1 score ($\frac{2\times\text{precision}\times\text{recall}}{\text{precision}+\text{recall}}$, where $\text{precision}=\frac{TP}{TP+FP}$ and $\text{recall}=\frac{TP}{TP+FN}$), we define TPTP as the number generated objects that are mapped to one (and only one) ground truth object, FPFP as the number of generated objects that are not mapped to any ground truth object, and FNFN as the number of ground truth objects that are not mapped to any generated object.
As for layout constraints, we first translate each generated constraint into some NL (i.e., an intuitive and human-understandable sentence). Then, we use Sentence-BERT to compute the embeddings of the translated generated constraints and ground truth NL-driven constraints. We then compute the scaled dot-product between every pair of generated and ground truth constraints, creating a confidence matrix (in some sense, a many-to-many mapping). For each entry in this matrix, if either (1) no object name from the corresponding ground truth constraint exists in the corresponding generated constraint or (2) it is smaller than a threshold $\tau_{l}$, the entry is zeroed out. Finally, to compute the F1 score, we define TPTP as the number of ground truth constraints that are mapped to at least one generated constraint, FPFP as the number of unmapped generated constraints, and FNFN as the number of unmapped ground truth constraints.
@@ 5.5.2. Stage II: Object-Query Coherence
For each acquired object, we (1) compute its smallest bounding sphere (with radius $r$ units), (2) place a camera $r\sin{\frac{FOV}{2}}$ units away from the object (where FOVFOV (in radian) is the field of view of the camera) pointing towards the front-facing side of the object, and (3) render the camera view using Blender on a white background. We then use BLIP-2 (Li et al., 2023a) (with its ITM head) and CLIP (Radford et al., 2021; Hessel et al., 2021) to measure the coherence between the formulated object query generated by Scenethesis and the rendered image. To reduce bias, in addition to the object query itself, we further pass a combination of “a 3D model of ” followed by the object query to the evaluation tools, and the final tool-wise score is the maximum between the two trials. The overall coherence is the arithmetic mean between the final BLIP and CLIP scores.